# Utilizing 5G NR SSB Blocks for Passive Detection and Localization of Low-Altitude Drones

### v1

@@ Appendix A Proof of proposition 1
$\begin{split}{\left[\text{FIM}(\Theta)\right]}_{d,d}&=2\text{Re}\biggl{\{}4\pi^{2}\frac{\rho\beta g^{2}}{M\sigma_{n}^{2}}\mathbb{E}\{\left\lvert\alpha\right\rvert^{2}\}\frac{f_{\Delta}^{2}}{c^{2}}L\sum_{n=1}^{N}n^{2}\biggl{\}}\\ &=\frac{4}{3}\pi^{2}\text{SNR}_{r}\frac{f_{\Delta}^{2}}{c^{2}}LN(N+1)(2N+1),\end{split}$ (13)
@@
where $[\text{FIM}(\Theta)]_{v,d}$$=$$[\text{FIM}(\Theta)]_{v,v}$, $\mathbb{E}\{\left\lvert\alpha\right\rvert^{2}\}$$=$$1$, $\text{SNR}_{r}$$=$$\frac{\rho\beta g^{2}}{\sigma_{n}^{2}M}$, $\sum_{n=1}^{N}n$$=$$\frac{N(N+1)}{2}$ ,$\sum_{n=1}^{N}n^{2}$$=$$\frac{N(N+1)(2N+1)}{6}$ and $T_{s}f_{\Delta}$$=$$1$. The Fishier information matrix is
@@
$\begin{split}{\text{CRB}(\Theta)}={\text{FIM}^{-1}(\Theta)}=\frac{1}{4\pi^{2}\text{SNR}_{r}NL\det({\text{FIM}(\Theta))}}\\ \begin{bmatrix}\frac{4}{3}\frac{T_{s}^{2}}{\lambda_{c}^{2}}(L+1)(2L+1)&\frac{1}{c\lambda_{c}}(L+1)(N+1)\\ \frac{1}{c\lambda_{c}}(L+1)(N+1)&\frac{1}{3}\frac{1}{c^{2}}f_{\Delta}^{2}(N+1)(2N+1)\end{bmatrix}\\ =\frac{1}{\pi^{2}\text{SNR}_{r}NL(7NL-N-L-5)}\\ \begin{bmatrix}3T_{s}^{2}c^{2}\frac{(2L+1)}{(N+1)}&\frac{9}{4}\lambda_{c}c\\ \frac{9}{4}\lambda_{c}c&\frac{3}{4}\lambda_{c}^{2}f_{\Delta}^{2}\frac{(2N+1)}{(L+1)}\end{bmatrix}.\end{split}$ (17)

### v2

@@ Appendix A Proof of proposition 1
$\begin{split}{\left[\text{FIM}(\Theta)\right]}_{d,d}&=2\text{Re}\biggl{\{}4\pi^{2}\frac{\rho\beta g^{2}}{M\sigma_{n}^{2}}\mathbb{E}\{\left\lvert\alpha\right\rvert^{2}\}\frac{f_{\Delta}^{2}}{c^{2}}L\sum_{n=1}^{N}n^{2}\biggl{\}}\\ &=\frac{4}{3}\pi^{2}\text{SNR}_{r}\frac{f_{\Delta}^{2}}{c^{2}}NL(N+1)(2N+1),\end{split}$ (13)
@@
where $[\text{FIM}(\Theta)]_{v,d}$$=$$[\text{FIM}(\Theta)]_{d,v}$, $\mathbb{E}\{\left\lvert\alpha\right\rvert^{2}\}$$=$$1$, $\text{SNR}_{r}$$=$$\frac{\rho\beta g^{2}}{\sigma_{n}^{2}M}$, $\sum_{n=1}^{N}n$$=$$\frac{N(N+1)}{2}$ ,$\sum_{n=1}^{N}n^{2}$$=$$\frac{N(N+1)(2N+1)}{6}$ and $T_{s}f_{\Delta}$$=$$1$. The Fishier information matrix is
@@
$\begin{split}{\text{CRB}(\Theta)}={\text{FIM}^{-1}(\Theta)}=\frac{1}{4\pi^{2}\text{SNR}_{r}NL\det({\text{FIM}(\Theta))}}\\ \begin{bmatrix}\frac{4}{3}\frac{T_{s}^{2}}{\lambda_{c}^{2}}(L+1)(2L+1)&\frac{1}{c\lambda_{c}}(L+1)(N+1)\\ \frac{1}{c\lambda_{c}}(L+1)(N+1)&\frac{1}{3}\frac{f_{\Delta}^{2}}{c^{2}}(N+1)(2N+1)\end{bmatrix}\\ =\frac{1}{\pi^{2}\text{SNR}_{r}NL(7NL-N-L-5)}\\ \begin{bmatrix}3T_{s}^{2}c^{2}\frac{(2L+1)}{(N+1)}&\frac{9}{4}\lambda_{c}c\\ \frac{9}{4}\lambda_{c}c&\frac{3}{4}\lambda_{c}^{2}f_{\Delta}^{2}\frac{(2N+1)}{(L+1)}\end{bmatrix}.\end{split}$ (17)